# Network-aided Efficient LLM Services With Denoising-inspired Prompt Compression

### v3

@@ 4.2 Double DQN Solution
To solve the complex JPPO optimization problem, we deploy a centralized Double DQN method to find optimal prompt compression and transmission strategies for $N$ users, as shown in Algorithm 1. The centralized control of the DDQN agent ensures fairness by dynamically allocating bandwidth and power based on real-time user demands and varying channel conditions. The key elements of the proposed Double DQN design are
@@
  • State: The state information includes the current fidelity of the transmitted message, SNR, and BEP. The state information of $n_{\rm th}$ user is captured in a 3-dimensional vector: $\left[\mathbf{f}_{n}(\eta_{n}),\gamma_{n},\right]$.
  • Action: The actions include selecting compression and power levels. The action space is denoted as $\mathcal{A}=\{\mathcal{A}_{1},\cdots,\mathcal{A}_{N}\}$, where $\mathcal{A}_{n}$ is the action of $n_{\rm th}$ user and consists of a tuple with discrete values of compression ratio and transmission power level. The compression ratio level is discrete values range from 00 to $4$. The transmission power level is discrete values ranging from 00 to $9$, which affects BEP.
  • Reward: The reward of the $n_{\rm th}$ user is $\mathcal{R}_{n}$, and in each episode, the agent accumulates rewards based on the actions. The reward function maximizes fidelity while minimizing penalties related to BEP and power usage.
@@
0:  Initialize the action space $\mathcal{A}=\{\mathcal{A}_{1},\cdots,\mathcal{A}_{N}\}$ and target privacy $(\epsilon,\delta)$.
0:  Reward.
1:  Initialization: $s_{0}$.
2:  for each episode $k\in\{1,...,K\}$ do
3:     Explore actions and obtain initial states.
4:     for each step in the episode do
5:        Take action $a_{t}$ according to current policy.
6:        Receive reward $r_{t}$ and observe the next state $s_{t+1}$.
7:        if $s^{\prime}$ is the final state then
8:           $y=\mathcal{R}^{\prime}$.
9:        else
10:           $y=\mathcal{R}^{\prime}+[\alpha r_{t+1}+\gamma\max_{a}Q(s_{t+1},a)-Q(s_{t},a_{t})]$.
11:        end if
12:        Sample a mini-batch of experiences from memory
13:        for each experience $(s,a,r,s^{\prime})$ in the batch: do
14:           Calculate target Q-values according to (13)
15:           Compute the loss according to (12)
16:           Perform gradient descent to minimize loss:
17:           Update $\epsilon$: $\epsilon=\max\left(\epsilon*\epsilon_{\text{decay}},\epsilon_{\min}\right)$
18:           Update the target Q-network every few episodes.
19:           Accumulate total reward and update state:
20:           $\text{Total reward}\>+=r$, $s=s^{\prime}$
21:        end for
22:     end for
23:  end for
24:  Update the policy parameter.
25:  Terminate the training when the policy converges or after a predefined number of iterations.
@@
$Q(s_{t},a_{t})\leftarrow Q(s_{t},a_{t})+\alpha r_{t+1}+\mu\max_{a}Q(s_{t+1},a)-Q(s_{t},a_{t}),$ (17)
where $Q(s_{t},a_{t})$ is the estimated Q-value updated by the Bellman equation for taking action $a_{t}$ in state $s_{t}$. $\alpha$ is the learning rate. $r_{t+1}$ is the reward received after taking action $a_{t}$ in state $s_{t}$ and transitioning to state $s_{t+1}$. $\mu$ is the discount factor. $\max_{a}Q(s_{t+1},a)$ is the maximum estimated Q-value for all possible actions in state $s_{t+1}$.
In DQN, the loss function is the Mean Squared Error (MSE) between the predicted Q-values and the target Q-values:
@@
where $L_{i}(\theta_{i})$ is the loss for the $i$-th iteration with parameter $\theta_{i}$. $r_{t+1}$ and $s_{t+1}$ are the reward and next state observed after taking action $a_{t}$ in state $s_{t}$. $Q_{\text{target}}(s_{t+1},a^{\prime};\theta_{i}^{*})$ is the target Q-value estimated by the target network with parameters $\theta_{i}^{*}$. $Q(s_{t},a_{t};\theta_{i})$ is the predicted Q-value by the current Q-network with parameters $\theta_{i}$.
The update steps of the Double DQN can be expressed as
@@ 5 Numerical Results
We design a customized environment with variable fidelity, SNR, and BEP to simulate the wireless network-aided LLM service framework. The centralized DQN agent manages the environment, which involves selecting compression and power levels. Its goal is to balance fidelity, minimize errors, and optimize power usage. The trained DRL model is designed with a modular architecture, making it adaptable to other types of networks and tasks. Its state representation, action selection, and reward mechanisms can be independently modified to accommodate communication networks. While the current implementation assumes a centralized optimization strategy, future work could explore decentralized and hybrid architectures to improve scalability and robustness. For instance, a distributed learning-based approach could allow multiple edge nodes to collaboratively optimize prompt compression and power allocation, thereby reducing the reliance on a single control unit. Moreover, integrating user-specific demands and heterogeneous network conditions into the optimization problem would allow for more adaptive and personalized service provisioning.
We use the LLMLingua platform, employing the SLM model with GPT-Neo 125M for prompt compression and GPT-J 6B to generate the response [35]. The simulations are carried out based on the MeetingBank-transcript dataset [36], where we select the long prompt data whose length exceeds $500$ tokens from the dataset to apply in our system simulations. Other parameter settings are listed in Table I.
TABLE I: Simulation parameter configuration
@@
$\alpha_{1}$, $\alpha_{2}$, and $\alpha_{3}$ $0.4$, $0.3$ and $0.3$
Total test runs range $[1,10]$
@@
  • Network Conditions: Our simulated MEC environment includes a total bandwidth of 20 MHz, a maximum transmission power of 10 dBm per user. Channel quality varies dynamically in the range of 0.1 to 1.
@@
Figure 4: The example illustrates wireless network-aided LLM services with SLM-based prompt compression, with a compression example of a single 16x compression ratio and an iterative 4 times 2x compression ratio.
[…]
Figure 5: One-shot reward values across different compression methods with respect to compression and power levels, tested using a single long prompt. Figs. 5 (a), (b), (c), and (d) show the reward performance of the baseline single-round compression method, the denoising-inspired prompt compression method with linear schedule, cosine schedule, and quadratic schedule, respectively.
In Fig. 5, we present the end-to-end performance of LLM inference in wireless network-aided LLM services after SLM compression and wireless transmission for a specific long prompt. Specifically, Fig. 5 (a) shows the results of single-step 16x compression, while Fig. 5 (b), Fig. 5 (c), and Fig. 5 (d) demonstrate the effects of different denoising-inspired prompt compression schedules (linear, cosine, and quadratic, respectively) under the same compression requirements. For each approach, the reward grid plot illustrates the relationship between compression levels (1-10) and power levels (1-10). We observe that the optimal rewards of the baseline compression method are achieved at moderate compression levels (2-5) to achieve high fidelity scores and moderate power levels (2-5) because of the energy constraint. Key insights from our analysis reveal an important trade-off between prompt compression ratio and transmit power. While higher compression ratios can reduce textural prompt volume and subsequent LLM inference time, allowing for relatively higher transmit power under energy constraints to reduce BEP, there is a risk of compromising performance due to loss of critical information affecting fidelity. Conversely, lower compression ratios maintain longer textural prompts but restrict transmit power settings, potentially increasing BEP and resulting in higher delays for LLM processing of longer prompts, leading to less noticeable performance gains. Another crucial insight is the variation in maximum achievable rewards across different denoising-inspired prompt compression schedules. For our test prompt, the optimal schedule demonstrates over $10\%$ improvement in reward value compared to single-step compression, which can translate to noticeably better inference results and reduced delay from the user’s perspective. This suggests that selecting an appropriate compression scheme based on the prompt’s context type after establishing a compression ratio requirement can be highly effective.
Figure 6: Performance metrics (i.e., reward and fidelity) of the DRL algorithm.
Fig. 6 demonstrates the training effects on long prompts exceeding 500 tokens from the MeetingBank-transcript dataset[36]. Specifically, Fig. 6 (a) illustrates the reward convergence patterns across different compression methods, showing how the Double DQN method effectively learns optimal policies for compression and power allocation. The cosine schedule demonstrates superior performance for meeting transcripts, as its smooth transitions (defined by $\sigma(t)=\frac{1-\cos(\pi t)}{2}$) are particularly well-suited for maintaining the coherence and context of conversational data. This effectiveness is especially notable for meeting transcripts, where the cosine schedule’s moderate compression steps at the beginning and end, with more aggressive compression in the middle stages, helps preserve both the meeting’s contextual flow and key discussion points. Fig. 6 (b) tracks the average fidelity across episodes, comparing different compression schedules and showing stabilization around $0.7$, which indicates robust information preservation even with significant compression. This high fidelity, combined with improved rewards, suggests our method successfully balances prompt compression levels and transmission power allocation, achieving efficient wireless network-aided LLM services while maintaining an acceptable average BEP.

### v4

@@ 4.2 Double DQN Solution
To solve the complex JPPO++ optimization problem, we deploy a centralized Double DQN method to find optimal prompt compression and transmission strategies for $N$ users, as shown in Algorithm 1.
4.2.1 Algorithm Design
The centralized control of the DDQN agent ensures fairness by dynamically allocating bandwidth and power based on real-time user demands and varying channel conditions. The key elements of the proposed Double DQN design are given as follows:
@@ 4.2.1 Algorithm Design
  • State: The state information includes the current fidelity of the transmitted message, SNR, and BEP. The state information of the $n_{\rm th}$ user is captured in a 3-dimensional vector: $\left[\mathbf{f}_{n}(\eta_{n}),\gamma_{n},\right]$.
  • Action: The actions include selecting compression and power levels. The action space is denoted as $\mathcal{A}=\{\mathcal{A}_{1},\ldots,\mathcal{A}_{N}\}$, where $\mathcal{A}_{n}$ is the action of the $n_{\rm th}$ user and consists of a tuple with discrete values of compression ratio and transmission power level. The compression ratio level is a discrete value range from 00 to $4$. The transmission power level is a discrete value ranging from 00 to $9$, which affects BEP.
  • Reward: The reward of the $n_{\rm th}$ user is $\mathcal{R}_{n}$. In each episode, the agent accumulates rewards based on its actions. The reward maximizes fidelity while minimizing penalties related to BEP and power usage.
@@
Input:
  • Action space $\mathcal{A}=\{\mathcal{A}_{1},\ldots,\mathcal{A}_{N}\}$
  • Target privacy parameters $(\epsilon,\delta)$
  • Learning rate $\xi$, exploration decay $\epsilon_{\text{decay}}$
Output:
  • Learned Q-network with optimized policy
1:Initialize Q-network and target Q-network with random weights
2:Initialize replay buffer and initial state $s_{0}$
3:for each episode $k\in\{1,\dots,K\}$ do
4:     Reset environment and observe initial state $s_{0}$
5:     while not episode terminated do
6:         Select action $a_{t}$ using $\epsilon$-greedy policy from Q-network
7:         Execute $a_{t}$ and observe reward $r_{t}$, and then go to next state $s_{t+1}$
8:         Store transition $(s_{t},a_{t},r_{t},s_{t+1})$ into replay buffer
9:         $s_{t}\leftarrow s_{t+1}$
10:         # Training step:
11:         Sample mini-batch of transitions $(s,a,r,s^{\prime})$ from replay buffer
12:         for each transition in the mini-batch do
13:              if $s^{\prime}$ is terminal then
14:                  $y\leftarrow r$
15:              else
16:                  Go to (22)
17:              end if
18:              Compute loss $\mathcal{L}=\left(y-Q\left(s,a;\theta\right)\right)^{2}$
19:              Perform gradient descent step on $\mathcal{L}$ to update Q-network
20:         end for
21:         Decay exploration rate: $\epsilon\leftarrow\max\left(\epsilon\cdot\epsilon_{\text{decay}},\epsilon_{\min}\right)$
22:         Periodically update target Q-network: $\theta^{-}\leftarrow\theta$
23:     end while
24:end for
@@
$Q\left(s_{t},a_{t}\right)\leftarrow Q\left(s_{t},a_{t}\right)+\xi r_{t+1}+\mu\max_{a}Q(s_{t+1},a)-Q\left(s_{t},a_{t}\right),$ (20)
where $Q\left(s_{t},a_{t}\right)$ is the estimated Q-value updated by the Bellman equation for taking action $a_{t}$ in state $s_{t}$. $\xi$ is the learning rate. $r_{t+1}$ is the reward received after taking action $a_{t}$ in state $s_{t}$ and transitioning to state $s_{t+1}$. $\mu$ is the discount factor. And the loss function is the Mean Squared Error (MSE) between the predicted Q-values and the target Q-values, as is given by:
@@
where $L_{i}(\theta_{i})$ is the loss for the $i$-th iteration with parameter $\theta_{i}$. $r_{t+1}$ and $s_{t+1}$ are the reward and next state observed after taking action $a_{t}$ in state $s_{t}$. $Q_{\text{target}}(s_{t+1},a^{\prime};\theta_{i}^{*})$ is the target Q-value estimated by the target network with parameter $\theta_{i}^{*}$. $Q(s_{t},a_{t};\theta_{i})$ is the predicted Q-value by current Q-network with parameter $\theta_{i}$.
Then the update of the Double DQN can be given by:
@@
4.2.2 Complexity Analysis
The computational complexity of the Double DQN algorithm is primarily determined by the number of training episodes, the size of the neural network, and the batch operations performed during training. Let $K$ denote the number of training episodes, $\mathcal{T}$ denote the average number of time steps per episode, $B$ denote the batch size, and $\mathcal{P}$ denote the number of parameters in the Q-network.
At each time step, the algorithm performs a forward pass through the Q-network for action selection, taking $\mathcal{O}(\mathcal{P})$ time. When updating the Q-network, a mini-batch of $B$ transitions is sampled from the replay buffer. For each sample, forward passes through both the current and target Q-networks, as well as a backward pass for optimization, are required. These operations contribute a per-step training cost of $\mathcal{O}(B\cdot\mathcal{P})$. Therefore, the total time complexity over all episodes is $\mathcal{O}(K\cdot\mathcal{T}\cdot B\cdot\mathcal{P})$.
In terms of space, the main contributors are the Q-network parameters, requiring $\mathcal{O}(\mathcal{P})$ memory, and the replay buffer, which stores up to $M$ transitions of size proportional to the state and action dimensions. Assuming each transition requires $\mathcal{O}(d)$ space, where $d$ is the dimension of the state vector, the total space complexity is $\mathcal{O}(\mathcal{P}+M\cdot d)$.
@@ 5 Numerical Results
We consider a practical simulation environment with variable fidelity, SNR, and BEP to support a wireless network-aided LLM service framework. The centralized DQN agent manages the environment, which involves selecting compression and power levels. Its goal is to balance fidelity, minimize errors, and optimize power usage. The trained DRL model is designed with a modular architecture, making it adaptable to other types of networks and tasks. Its state representation, action selection, and reward mechanisms can be independently modified to accommodate communication networks.
5.1 Experimental Setup
Following the LLMLingua platform, we employ the SLM, i.e., GPT-Neo 125M, for prompt compression and the LLM, i.e., GPT-J 6B, to generate responses for users [35]. The simulations are carried out based on the MeetingBank-transcript dataset [36] and LongBench datset  [37], where we select the long prompt data whose length exceeds $500$ tokens from the dataset to apply in our system simulations. Other parameter settings are in Table II.
TABLE II: Simulation Parameter Configuration
@@ 5.1 Experimental Setup
$\alpha_{1}$, $\alpha_{2}$, $\alpha_{3}$ $0.4$, $0.3$, $0.3$
Total Test Runs Range $\left[1,10\right]$
@@
  • Network Conditions: Our simulated MEC environment includes a total bandwidth of $20$ MHz, a maximum transmission power of $10$ dBm per user. In the MEC environment, the transmission power ranges from a minimum of 1 dBm (approximately 0.00126 W) to a maximum of 10 dBm (0.01 W). We divide the range of power into 10 power levels evenly, where each level increases by 1 dBm, corresponding to a gradual increase in power from low to high. Additionally, the channel quality of the network varies dynamically in the range of $0.1$ to $1$.
@@
Figure 4: The example illustrates wireless network-aided LLM services with SLM-based prompt compression, with a compression example of a single 16×16$\times$16 × compression ratio and an iterative $4$ times 2×2$\times$2 × compression ratio.
5.2 Experiments Performance Analysis
To demonstrate the effectiveness and generalizability of our proposed JPPO++ framework, we design a sequence of experiments to answer the following research questions:
  • Q1. Effectiveness: Does denoising-inspired prompt compression improve fidelity compared to one-shot compression?
  • Q2. Necessity: How do compression ratio and transmission power jointly affect QoS in network-aided LLM services?
  • Q3. Adaptivity: Can DRL effectively optimize prompt compression and power control under dynamic network and task conditions?
5.2.1 Q1. Effectiveness
Fig. 4 compares prompt compression using a single-round 16×16$\times$16 × compression ratio with an iterative method applying 2×2$\times$2 × compression over 4 times. Based on the fidelity metric $\mathbf{f}$, the iterative strategy consistently outperforms the single-step counterpart across all three components, demonstrating better semantic preservation and communication robustness. The iterative method achieves a final fidelity score of $0.27$, compared to $0.18$ for the single-round compression. The total processing time is $5.90$ seconds, including $2.92$ seconds for compression, $2.48$ seconds for LLM response generation, and $0.50$ seconds for transmission. From the DRL perspective, it yields a higher reward score of $26.40$ while consuming $8.18$ units of power.
In contrast, the single-round 16×16$\times$16 × compression approach consumes less power (i.e., $2.90$ units), but produces a lower reward of $17.74$ and a longer total processing time of $11.25$ seconds that is nearly double that of the iterative method. This includes $6.59$ seconds for compression, $4.12$ seconds for response generation, and $0.55$ seconds for transmission. These results confirm that denoising-inspired iterative compression in JPPO++ not only improves fidelity but also significantly reduces end-to-end latency, making it a more efficient strategy for LLM prompt compression under wireless constraints. We further validate the effectiveness of our framework using wider tests in the MeetingBank-transcript dataset [36], where prompts exceed $500$ tokens. When prioritizing aggressive compression while tolerating up to a $30\%$ fidelity drop, JPPO++ supports compression ratios up to 16×16$\times$16 ×. Under this setting, service time is significantly reduced. The single-round baseline achieves a $42.3\%$ reduction in total service time compared to the no-compression case, while our iterative denoising-based method improves it further to $46.5\%$.
5.2.2 Q2. Necessity
TABLE III: One-shot reward values of the baseline single-round compression method.
Power Level
1 2 3 4 5 6 7 8 9 10
Compression 1 10.11 10.64 10.76 10.71 10.58 10.39 10.18 9.95 9.71 9.45
2 11.12 11.66 11.77 11.72 11.59 11.41 11.20 10.97 10.72 10.47
3 11.40 11.93 12.05 12.00 11.87 11.69 11.47 11.24 11.00 10.75
4 11.41 11.94 12.06 12.01 11.87 11.69 11.48 11.25 11.01 10.75
5 11.31 11.85 11.96 11.91 11.78 11.60 11.39 11.16 10.91 10.66
6 11.19 11.72 11.84 11.79 11.65 11.47 11.26 11.03 10.79 10.53
7 11.04 11.58 11.69 11.64 11.51 11.33 11.12 10.89 10.64 10.39
8 10.90 11.43 11.55 11.50 11.36 11.18 10.97 10.74 10.50 10.24
9 10.75 11.28 11.40 11.35 11.21 11.03 10.82 10.59 10.35 10.09
10 10.60 11.13 11.25 11.20 11.07 10.88 10.67 10.44 10.20 9.94
TABLE IV: One-shot reward values of the denoising-inspired prompt compression method with linear schedule.
Power Level
1 2 3 4 5 6 7 8 9 10
Compression 1 10.25 10.93 11.17 11.25 11.24 11.17 11.07 10.94 10.79 10.63
2 11.27 11.94 12.18 12.26 12.24 12.18 12.07 11.94 11.79 11.63
3 11.54 12.21 12.46 12.53 12.51 12.44 12.34 12.21 12.06 11.89
4 11.55 12.21 12.46 12.53 12.51 12.44 12.34 12.20 12.05 11.88
5 11.45 12.12 12.36 12.43 12.41 12.34 12.23 12.10 11.95 11.78
6 11.32 11.99 12.23 12.30 12.28 12.20 12.09 11.96 11.80 11.63
7 11.18 11.84 12.08 12.15 12.13 12.05 11.94 11.80 11.65 11.47
8 11.03 11.69 11.93 12.00 11.97 11.90 11.78 11.65 11.49 11.31
9 10.88 11.54 11.78 11.84 11.82 11.74 11.62 11.48 11.32 11.15
10 10.73 11.39 11.62 11.69 11.66 11.58 11.46 11.32 11.16 10.98
TABLE V: One-shot reward values of the denoising-inspired prompt compression method with cosine schedule.
Power Level
1 2 3 4 5 6 7 8 9 10
Compression 1 9.96 10.50 10.62 10.57 10.44 10.27 10.06 9.83 9.59 9.34
2 10.84 11.38 11.50 11.45 11.33 11.15 10.95 10.72 10.48 10.24
3 10.98 11.52 11.64 11.60 11.48 11.31 11.10 10.88 10.65 10.40
4 10.85 11.39 11.52 11.48 11.36 11.19 10.99 10.78 10.54 10.30
5 10.62 11.17 11.30 11.27 11.15 10.98 10.79 10.57 10.34 10.10
6 10.36 10.92 11.05 11.02 10.90 10.74 10.55 10.33 10.11 9.87
7 10.10 10.65 10.79 10.76 10.65 10.49 10.30 10.09 9.86 9.63
8 9.83 10.38 10.52 10.50 10.39 10.23 10.04 9.84 9.62 9.39
9 9.56 10.12 10.26 10.24 10.13 9.98 9.79 9.59 9.37 9.14
10 9.29 9.85 10.00 9.98 9.88 9.72 9.54 9.34 9.13 8.90
TABLE VI: One-shot reward values of the denoising-inspired prompt compression method with quadratic schedule.
Power Level
1 2 3 4 5 6 7 8 9 10
Compression 1 4.44 8.61 10.29 10.80 10.68 10.22 9.56 8.80 7.98 7.15
2 5.62 9.92 11.77 12.45 12.51 12.22 11.72 11.10 10.43 9.72
3 6.01 10.36 12.27 13.04 13.18 12.97 12.55 12.01 11.40 10.76
4 6.10 10.46 12.41 13.22 13.41 13.25 12.88 12.38 11.82 11.21
5 6.09 10.44 12.40 13.24 13.46 13.33 12.99 12.52 11.99 11.41
6 6.03 10.36 12.33 13.18 13.42 13.31 12.99 12.55 12.03 11.48
7 5.96 10.26 12.23 13.08 13.34 13.24 12.94 12.52 12.02 11.48
8 5.87 10.15 12.11 12.97 13.23 13.15 12.86 12.45 11.96 11.43
9 5.79 10.03 11.98 12.84 13.11 13.03 12.75 12.35 11.88 11.36
10 5.70 9.91 11.85 12.71 12.98 12.91 12.64 12.24 11.77 11.27
Tables III-VI show the end-to-end DRL reward values of wireless network-aided LLM service under different combinations of compression levels and transmission power levels. Specifically, Table III presents the results for baseline single-round 16×16$\times$16 × compression, while Tables IV-VI show the outcomes of denoising-inspired iterative compression with linear, cosine, and quadratic scheduling strategies, respectively. For each approach, the reward matrix reflects the impact of varying compression ratios (i.e., rows) and transmission power levels (i.e., columns) on final system performance.
We observe that, across all methods, the highest reward values are typically achieved at moderate compression levels (e.g., 2×2$\times$2 ×-5×5$\times$5 ×) and moderate power levels (e.g., 2-5). This pattern reflects a fundamental trade-off shaped by energy constraints and semantic preservation. On the one hand, aggressive compression reduces prompt length and transmission overhead, enabling higher transmit power within the same energy budget. However, it also increases the risk of discarding critical semantic content, which in turn degrades the quality of LLM inference. On the other hand, lower compression ratios preserve more of the original prompt, thereby improving fidelity, but result in longer input sequences. This increases LLM inference latency and limits the available transmission power, potentially raising the BEP and reducing the overall reward. Similarly, power levels exhibit a non-linear effect: very low power leads to high BEP and unreliable transmission, while very high power may violate energy constraints. The observed reward surface, therefore, captures the complex interplay between compression aggressiveness and transmission reliability, highlighting the necessity of joint optimization.
Therefore, a key insight from this analysis is that compression ratio and power control are not independently tunable. Thus, joint optimization is necessary to strike a context-aware balance. Another critical observation lies in the differences among scheduling strategies. The denoising-inspired schemes with dynamic compression progression (particularly the quadratic and cosine schedules) achieve higher peak rewards compared to the single-shot baseline. For instance, the best-performing configuration under the quadratic schedule exceeds the baseline reward by over $10\%$. This indicates that gradually compressing prompts in stages, especially with curvature-adaptive schedules, better preserves key information and aligns more effectively with LLM inference patterns. These findings suggest that selecting a compression trajectory tailored to the prompt structure, according to real scenarios, can provide substantial performance benefits.
5.2.3 Q3. Adaptivity
Figure 5: The convergence performance of reward over 100 iterations for the DRL algorithm.
Fig. 5 presents the reward convergence curves during the training process over $100$ training episodes using our Double DQN-based solution for the JPPO++ framework on long prompts from the MeetingBank-transcript dataset [36]. The DRL agent successfully learns optimal compression and power control policies under different iterative schedules. Notably, the cosine schedule consistently achieves the highest reward, demonstrating its effectiveness in preserving semantic coherence in long, structured dialogues. Its smooth transition function $\sigma(t)=\frac{1-\cos(\pi t)}{2}$ enables gradual compression, with moderate intensity at the beginning and end and stronger compression in the middle. This pattern is particularly suited for meeting transcripts, where key information is interleaved with contextual background. Compared to the static baseline, the learned policy yields faster convergence and higher final performance, confirming that JPPO++ can adaptively balance fidelity and efficiency in real-time through policy learning.
(a) Transmission time reduction percentage
(b) Service time reduction percentage
Figure 6: The illustration of transmission time and service time reduction percentages of using the denoising-inspired prompt compression method in JPPO++ compared to the single-round compression baseline.
As part of the testing phase, Fig. 6 presents the transmission time and service time reduction for five randomly selected samples from the MeetingBank-transcript dataset. Specifically, transmission time reduction corresponds to the time reduced by the denoising-inspired prompt compression method in JPPO++ compared to the single-round compression baseline. Service time, as illustrated in Fig. 1, is defined as the sum of compression time, transmission time, and LLM inference time, capturing the end-to-end efficiency of the mobile LLM service. Across all samples, the denoising-inspired method consistently outperforms the baseline, achieving transmission time reductions of approximately $10\%$ to $65\%$ and service time reductions of $39\%$ to $52\%$. These results confirm that the proposed JPPO++ enhances data delivery efficiency while preserving task quality. The consistent improvements across diverse prompts and varying content structures demonstrate the robustness and suitability for real-world mobile networks.
(a) Multi-news category
(b) GovReport category